OpenStudy (fairy733189):

Which graph shows the quadratic function y = 3x^2 - 12x + 10?

OpenStudy (zzr0ck3r):

what is the vertex of the quadratic?

OpenStudy (fairy733189):

hang on let me put the graphs

OpenStudy (zzr0ck3r):

we probably dont need them

OpenStudy (zzr0ck3r):

do you know what the vertex is and how to find it?

OpenStudy (fairy733189):

yes, I know that in order to find the vertex you have to do -b/2a

OpenStudy (zzr0ck3r):

well that will tell you the x value, so do that first

OpenStudy (fairy733189):

ok: -(-12)/2(32) = 12/ 64 = 3/16

OpenStudy (zzr0ck3r):

where did you get 32?

OpenStudy (fairy733189):

isnt that the a value?

OpenStudy (fairy733189):

oh sorry I thought it was 32 I just realized it was 3

OpenStudy (zzr0ck3r):

\(ax^2+bx+c\) 3x^2 - 12x + 10

OpenStudy (zzr0ck3r):

correct, 3

OpenStudy (fairy733189):

so it would be: -(-12)/2(3) = 12/6 = 2

OpenStudy (zzr0ck3r):

ok, now plug 2 back into the function and what do you get?

OpenStudy (fairy733189):

y = 3(2)^2 - 12(2) + 10 y = 12 - 24 + 10 y = -12 + 10 I would get y = -2

OpenStudy (zzr0ck3r):

ok so the vertex is \((2,-2)\) so which graph is it?

OpenStudy (fairy733189):

would it be graph d?

OpenStudy (zzr0ck3r):

correct

OpenStudy (fairy733189):

Thank you so much!!! Do you think you can help me with another problem?

OpenStudy (zzr0ck3r):

sure close this and ask a new one

OpenStudy (fairy733189):

ok thanks!